OpenStudy (anonymous):

Quadrilateral ABCD is a rhombus. Its diagonals AC and BD intersects at E. If AE = 3x + 3 and CE = 5x-3 , find AC.

OpenStudy (anonymous):

Diagonals of a rhombus bisect each other, therefore, AE = CE 3x + 3 = 5x - 3 2x = 6 x = 3 So, AC = AE + EC = 12 + 12 = 24 Nore that when x = 3, 3x + 3 = 12

OpenStudy (anonymous):

Thank you. :) How about. If DE=2x + 2 and BE=3x-1 , find BD

OpenStudy (anonymous):

Exact same question as before...DE = BE...now you continue just like I did in the last problem